OpenStudy (anonymous):

Fully factorize X^3-X^2-6x

OpenStudy (campbell_st):

well factor out an x and you get \[x(x^2 -x - 6)\] now factor the quadratic expression

OpenStudy (anonymous):

Thank you so much! Just got a curve equation and its awful XD

OpenStudy (anonymous):

Screw it will ask it here

OpenStudy (anonymous):

The point P with the coordinates (3,m) lies on the curve y=x^2 + kx. At P the gradient of the curve is 8. Find the values of the constants k and m.

OpenStudy (campbell_st):

if the point is on the curve then you knoa by substituting x = 3 that y = 3^2 + 3k so y = 9 + 3k or using the point m = 9 + 3k (Equation 1) now take the derivative of the curve y' = 2x + k you know the slope is 8... so y'=8 when x = 3 substitute them to find k 8 = 6 + k then so when you get k, go back and substitute it into equation 1 to find m. hope it helps